OpenStudy (anonymous):

find the vertex of the parabola: f(x)= x^2-x+1

OpenStudy (dumbcow):

a=1 b=-1 c=1 x = -b/2a x= -(-1)/2(1) = 1/2 y = (1/2)^2 - 1/2 +1 y = 1/4 - 1/2 +1 y = 3/4 vertex = (1/2, 3/4) Also you could use completing the square to put it in vertex form y = a(x-h)^2 +k where vertex = (h,k) take half of b b/2 = -1/2 this will be the "-h" part now square it and subtract from c (-1/2)^2 = 1/4 1 - 1/4 = 3/4 y = (x- 1/2)^2 + 3/4 vertex = (1/2, 3/4)